# TensorHyper-VQC: A Tensor-Train-Guided Hypernetwork for Robust and Scalable Variational Quantum Computing

### v2

@@ Abstract
Variational Quantum Computing (VQC) faces fundamental scalability barriers, primarily due to the presence of barren plateaus and its sensitivity to quantum noise. To address these challenges, we introduce TensorHyper-VQC, a novel tensor-train (TT)-guided hypernetwork framework that significantly improves the robustness and scalability of VQC. Our framework fully delegates the generation of quantum circuit parameters to a classical TT network, effectively decoupling optimization from quantum hardware. This innovative parameterization mitigates gradient vanishing, enhances noise resilience through structured low-rank representations, and facilitates efficient gradient propagation. Grounded in Neural Tangent Kernel and statistical learning theory, our rigorous theoretical analyses establish strong guarantees on approximation capability, optimization stability, and generalization performance. Extensive empirical results across quantum dot classification, Max-Cut optimization, and molecular quantum simulation tasks demonstrate that TensorHyper-VQC consistently achieves superior performance and robust noise tolerance, including hardware-level validation on a 156-qubit IBM Heron processor. These results position TensorHyper-VQC as a scalable and noise-resilient framework for advancing practical quantum machine learning on near-term devices.
@@ 1 Introduction
Variational Quantum Computing (VQC) [cerezo2021variational] has emerged as a prominent paradigm for exploiting the computational capabilities of Noisy Intermediate-Scale Quantum (NISQ) devices [preskill2018quantum]. VQC integrates quantum circuits and classical optimization routines to address computationally challenging problems across various domains [biamonte2017quantum, cerezo2022challenges, caro2022generalization, mcclean2016theory, power_data, schuld2019quantum]. Despite its significant promise, the practical realization of VQC’s potential is substantially hindered by fundamental obstacles related to its trainability, robustness against quantum noise, and overall scalability as problem sizes increase [holmes2022connecting]. A primary impediment to scalable VQC is the barren plateau phenomenon [larocca2025barren, mcclean2018barren, martin2023barren], where gradients vanish exponentially with increasing qubit count or circuit depth, significantly degrading result fidelity and destabilizing training [beer2020training, sharma2022trainability, pesah2021absence].
@@
To overcome these critical limitations, as shown in Fig. 1, we introduce TensorHyper-VQC, a tensor-train (TT)-guided hypernetwork approach designed to facilitate scalable and robust VQC. The central innovation of TensorHyper-VQC lies in its parameterization strategy. Instead of directly optimizing parameters on the quantum device, a classical TT network [oseledets2011tensor, yang2017tensor, qi2023exploiting] is trained to generate the parameters for the VQC. Consequently, the entire optimization loop, including gradient computation and parameter updates, is offloaded to the classical TT-cores. In this architecture, the quantum circuit functions solely as a fixed forward-pass evaluator, executing with the parameters supplied by the TT network to provide the objective function value. This reparameterization strategy enables a clear separation between classical optimization and quantum inference, ensuring consistent gradient propagation even under realistic hardware noise.
@@
Figure 1: An illustration of the TensorHyper-VQC framework. A classical TT network generates the parameters $\hat{\textbf{w}}=[\alpha^{(1)}_{1:U},\beta^{(1)}_{1:U},\gamma^{(1)}_{1:U},...,\alpha^{(L)}_{1:U},\beta^{(L)}_{1:U},\gamma^{(L)}_{1:U}$$^{\top}$ that define the Pauli rotation gates $R_{X}(\alpha^{(l)}_{u}),R_{Y}(\beta^{(l)}_{u}),R_{Z}(\gamma^{(l)}_{u})$ (where $u\in[U],l\in[L]$) in a VQC with $L$ depths and $U$ qubits. The TT-generated parameters are injected into the fixed quantum circuit, which operates solely in inference mode. The loss is computed from quantum measurements, and gradients are backpropagated only through the classical TT-cores. Gaussian noise is injected during training as the input to the TT network. This architecture decouples optimization from quantum hardware, mitigating barren plateaus and improving noise resilience.
@@
This hypernetwork-based parameterization fundamentally reshapes the VQC optimization landscape, offering distinct advantages in terms of scalability and robustness. By confining the optimization process to the classical TT network, TensorHyper-VQC significantly mitigates the barren plateau problem, as gradients are backpropagated through the well-conditioned, deterministic structure of the TT network rather than the potentially deep and complex quantum circuit. The framework’s resilience to quantum noise is also markedly enhanced, as the trainable parameters reside entirely within the classical domain, insulating gradient updates from the stochasticity of quantum measurements and hardware imperfections. Moreover, the low-rank structure inherent in the TT representation further contributes to noise suppression by effectively averaging measurement noise across the TT-core gradients, a phenomenon theoretically linked to variance reduction. Our Neural Tangent Kernel (NTK) [bietti2019inductive, jacot2018neural, liu2022representation] analysis supports the improved trainability afforded by this structured parameterization.
@@
Our theoretical investigations confirm the enhanced representation capability, trainability, and generalization properties of TensorHyper-VQC. Experimentally, TensorHyper-VQC demonstrates superior performance and robustness across diverse applications, including quantum machine learning (e.g., quantum dot classification [kalantre2019machine, czischek2021miniaturizing]), combinatorial optimization (e.g., Max-Cut [zhou2020quantum, zhu2022adaptive]), and quantum simulation (e.g., LiH Hamiltonians [o2016scalable, georgescu2014quantum]). In all cases, TensorHyper-VQC outperforms conventional VQC and exhibits strong noise resilience across various noise models. These findings establish TensorHyper-VQC as a powerful and versatile computational framework that advances state-of-the-art variational quantum computing.
@@ 2.1 The TensorHyper-VQC Framework
Figure 2: The architecture of variational quantum computing. The classical input data are encoded into quantum states $|x_{1}\rangle$, $|x_{2}\rangle$, …, $|x_{U}\rangle$ via data encoding (e.g., angle encoding and amplitude encoding). The shaded region denotes $L$ repeated parameterized quantum circuit block, consisting of trainable rotation gates $R_{X}(\alpha^{(l)}_{u})$, $R_{Y}(\beta^{(l)}_{u})$, $R_{Z}(\gamma^{(l)}_{u})$, and entangling layers across quantum channels. These blocks are stacked to increase circuit depth and expressive capacity. Finally, the circuit is measured in the Pauli-Z basis, yielding expectation values $\langle\sigma_{z}^{(1)}\rangle$, $\langle\sigma_{z}^{(2)}\rangle$, …, $\langle\sigma_{z}^{(U)}\rangle$ that serve as outputs for optimizaiton.
It is important to note that throughout this work, we employ a single, consistent VQC architecture across all experiments and applications. The parameterized quantum circuit block, illustrated in Figure 2, is fixed in structure: it consists of alternating layers of single-qubit rotations $R_{X}(\alpha^{(l)}_{u})$, $R_{Y}(\beta^{(l)}_{u})$, $R_{Z}(\gamma^{(l)}_{u})$ and entangling gates across $U$ qubits, stacked to achieve the desired circuit depth $L$. By holding the VQC design constant, our results isolate the contribution of the TensorHyper-VQC parameterization strategy, rather than relying on task-specific ansatz engineering. This highlights that the observed gains in trainability, robustness, and scalability stem directly from the TT-based hypernetwork, demonstrating its versatility as a general-purpose parameter generator for VQC.
@@
Finally, a consistent TT hypernetwork architecture is adopted across all applications. The TT ranks and core dimensions are chosen to balance expressivity and generalization, but the same generator design is applied across tasks throughout. This uniformity underscores that the improvements in scalability, robustness, and trainability derive from the TT-guided parameterization principle itself rather than from task-specific quantum architectural search.
@@ 2.2 Theoretical Guarantees
$\mathcal{T}_{\rm tv}=\nabla_{\boldsymbol{\theta}}f_{\boldsymbol{\theta}}(\textbf{x})\nabla_{\boldsymbol{\theta}}f_{\boldsymbol{\theta}}(\textbf{x})^{\top}=\left(\frac{\partial f_{\boldsymbol{\theta}}}{\partial\hat{\textbf{w}}}\cdot\frac{\partial\hat{\textbf{w}}}{\partial\boldsymbol{\theta}}\right)\left(\frac{\partial f_{\boldsymbol{\theta}}}{\partial\hat{\textbf{w}}}\cdot\frac{\partial\hat{\textbf{w}}}{\partial\boldsymbol{\theta}}\right)^{\top}.$ (8)
@@
$\lambda_{\rm min}(\mathcal{T}_{\rm tv})=\lambda_{\rm min}(\textbf{J}_{\hat{\textbf{w}}}\textbf{J}_{\hat{\textbf{w}}}^{\top}).$ (11)
@@
This shows that the effective NTK eigenvalue is not smaller, and under mild dispersion assumptions, can be strictly larger, ensuring faster convergence and mitigating vanishing gradients. Accordingly, at epoch $t$, the optimization error satisfies:
@@
$\frac{\partial\hat{\mathcal{R}}}{\partial\mathcal{G}_{k}}=\left(\frac{\partial\hat{\mathcal{R}}}{\partial\hat{\textbf{w}}}+\boldsymbol{\tau}\right)\cdot\frac{\partial\hat{\textbf{w}}}{\partial\mathcal{G}_{k}}=\sum\limits_{u=1}^{3UL}\frac{\partial\hat{\mathcal{R}}}{\partial\hat{w}_{u}}\cdot\frac{\partial\hat{w}_{u}}{\partial\mathcal{G}_{k}}+\sum\limits_{u=1}^{3UL}\tau_{u}\frac{\partial\hat{w}_{u}}{\partial\mathcal{G}_{k}}.$ (18)
@@ 2.4 Application to Quantum Machine Learning: Quantum Dot Classification
We first apply TensorHyper-VQC to a binary classification task on charge-stability diagrams from semiconductor quantum dots [tang2015storage, kalantre2019machine, gualtieri2025qdsim] to validate our theoretical framework. The objective is to classify diagrams into single-dot (Label 0) or double-dot (Label 1) systems, a crucial task for characterizing and controlling quantum devices. The experimental dataset consists of $50\times 50$-pixel images associated with the quantum dot diagrams, including $2,000$ diagrams generated under realistic device-level noise conditions. We randomly partition the diagrams into $1,800$ training and $200$ test samples to evaluate the models’ generalization performance.
@@
The proposed TensorHyper-VQC architecture integrates a classical TT network as a parameter generator for VQC. The TT network is configured with input dimensions $[5,10,5,10]$, output dimensions $[4,2,3,9]$, and TT-ranks $[1,2,2,2,1]$, resulting in a compact representation with only $576$ trainable parameters (including bias). The TT network maps vectorized quantum-dot features to all rotation angles required by the VQC, effectively parameterizing $20$ qubits across $6$ variational layers (360 quantum gate parameters). All models are trained using the cross-entropy loss [mao2023cross] based on the Adam optimizer [adam:2014] with a fixed learning rate $3\times 10^{-2}$. We use a fixed random seed across all experiments to ensure reproducibility and fairness, and we initialize parameters from a normal distribution with mean 0 and variance 1.
To comprehensively benchmark performance, we compare TensorHyper-VQC against multiple baselines, including (i) a standard VQC with randomly initialized parameters, (ii) MLPHyper-VQC, where a multilayer perceptron replaces the TT network as the parameter generator, (iii) TTN+VQC, a hybrid model combining a TT encoder and VQC [qi2023theoretical, chen2021end], and (iv) ResNet18 [szegedy2017inception] and ResNet50 [wen2020transfer] enhanced with low-rank adaptation (LoRA) [hu2022lora, liu2024dora] for classical comparison. Furthermore, we conduct robustness analyses by injecting a variety of realistic noise models into the quantum circuits, including depolarizing, dephasing, two-Pauli, and readout noise, with error probabilities ranging from $0.1\%$ to $0.3\%$, simulating typical NISQ-era conditions.
To further assess resilience to noise, we incorporate standard barren-plateau mitigation and error-reduction techniques, including layer-wise training (LWT) [skolik2021layerwise], which progressively optimizes circuit depth to maintain gradient signal, and local cost function (LCF) design [cerezo2021cost], which mitigates gradient vanishing by focusing on local observables. In addition, we employ standard quantum error mitigation approaches such as zero-noise extrapolation (ZNE) [he2020zero], readout error mitigation (REM) [smith2021qubit], and their combination (REM+ZNE). We evaluate the performance of all these strategies on both simulated and physical backends. Finally, we validate our approach on a 156-qubit IBM Quantum processor (Heron r2), demonstrating the deployability of TensorHyper-VQC on real hardware and its robustness against device-level imperfections, such as decoherence, crosstalk, and measurement errors.
@@
Robustness Under Multi-Noise Conditions. We next evaluate the robustness under realistic quantum noise models [cross2019validating, resch2021benchmarking], including depolarizing, dephasing, two-Pauli, and readout noise channels, as illustrated in Fig. 3(c). Even as noise levels as high as $0.3\%$, TensorHyper-VQC maintains strong performance ($\geq 85\%$), significantly outperforming noise-agnostic VQC baselines. This robustness arises from the TT-based parameterization, which smooths the optimization landscape and enhances noise tolerance. Furthermore, as shown in Fig. 3(d), TenosrHyper-VQC exhibits graceful degradation when scaling from $20$ qubits down to $16$ or $12$ qubits, demonstrating robustness to hardware resource constraints while retaining competitive performance.
Comparison with Quantum Error Mitigation. We also compare TensorHyper-VQC with standard quantum error mitigation techniques, including ZNE, REM, and their combination as shown in Fig. 3(e). Remarkably, TensorHyper-VQC achieves a final accuracy of $94.0\%$, outperforming VQC enhanced with ZNE+REM ($55.7\%$) by nearly $40$ percentage points. This highlights that our tensor-structured parameterization intrinsically mitigates the effects of noise by reducing parameter variance and constraining quantum state evolution, often outperforming post-hoc mitigation methods that require additional circuit executions and calibration overhead.
@@
Table 1: Ablation study on TT-rank selection under the TensorHyper-VQC’s configuration and noisy quantum setting described in Fig. (e).

### v3

@@ Abstract
Variational Quantum Computing (VQC) faces fundamental scalability barriers, primarily due to barren plateaus and sensitivity to quantum noise. To address these challenges, we introduce TensorHyper-VQC, a novel tensor-train (TT)-guided hypernetwork framework that significantly improves the robustness and scalability of VQC. Our framework fully delegates the generation of quantum circuit parameters to a classical TT network, effectively decoupling optimization from quantum hardware. This innovative parameterization mitigates gradient vanishing, enhances noise resilience through structured low-rank representations, and facilitates efficient gradient propagation. Grounded in Neural Tangent Kernel and statistical learning theory, our rigorous theoretical analyses establish strong guarantees on approximation capability, optimization stability, and generalization performance. Extensive empirical results across quantum dot classification, Max-Cut optimization, and molecular quantum simulation tasks demonstrate that TensorHyper-VQC consistently achieves superior performance and robust noise tolerance, including hardware-level validation on a 156-qubit IBM Heron processor. These results position TensorHyper-VQC as a scalable and noise-resilient framework for advancing practical quantum machine learning on near-term devices.
@@ 1 Introduction
Variational Quantum Computing (VQC) [6] has emerged as a prominent paradigm for exploiting the computational capabilities of Noisy Intermediate-Scale Quantum (NISQ) devices [34]. VQC integrates quantum circuits and classical optimization routines to address computationally challenging problems across various domains [2, 5, 4, 30, 18, 39]. Despite its significant promise, the practical realization of VQC’s potential is substantially hindered by fundamental obstacles related to its trainability, robustness against quantum noise, and overall scalability as problem sizes increase [16]. A primary impediment to scalable VQC is the barren plateau phenomenon [24, 29, 28], where gradients vanish exponentially with increasing qubit count or circuit depth, significantly degrading result fidelity and destabilizing training [1, 40, 33].
@@
To overcome these critical limitations, as shown in Fig. 1, we introduce TensorHyper-VQC, a tensor-train (TT)-guided hypernetwork approach designed to facilitate scalable and robust VQC. The central innovation of TensorHyper-VQC lies in its parameterization strategy. Instead of directly optimizing parameters on the quantum device, a classical TT network [32, 47, 36] is trained to generate the parameters for the VQC. Consequently, the entire optimization loop, including gradient computation and parameter updates, is offloaded to the classical TT-cores. In this architecture, the quantum circuit serves solely as a fixed forward-pass evaluator, taking the parameters supplied by the TT network to compute the objective function value. This reparameterization strategy enables a clear separation between classical optimization and quantum inference, ensuring consistent gradient propagation even in the presence of realistic hardware noise.
@@
Figure 1: An illustration of the TensorHyper-VQC framework. A classical TT network generates the parameters $\hat{\textbf{w}}=[\alpha^{(1)}_{1:U},\beta^{(1)}_{1:U},\gamma^{(1)}_{1:U},...,\alpha^{(L)}_{1:U},\beta^{(L)}_{1:U},\gamma^{(L)}_{1:U}$$^{\top}$ that define the Pauli rotation gates $R_{X}(\alpha^{(l)}_{u}),R_{Y}(\beta^{(l)}_{u}),R_{Z}(\gamma^{(l)}_{u})$ (where $u\in[U],l\in[L]$) in a VQC with $L$ depths and $U$ qubits. The $T$-generated parameters are injected into the fixed quantum circuit, which operates solely in inference mode. The loss is computed from quantum measurements, and gradients are backpropagated only through the classical TT-cores. Gaussian noise is injected during training as the input to the TT network. This architecture decouples optimization from quantum hardware, mitigating barren plateaus and improving noise resilience.
@@
This hypernetwork-based parameterization fundamentally reshapes the VQC optimization landscape, offering distinct advantages in terms of scalability and robustness. By confining the optimization to the classical TT network, TensorHyper-VQC significantly mitigates the barren-plateau problem, as gradients are backpropagated through the well-conditioned, deterministic structure of the TT network rather than through the potentially deep and complex quantum circuit. The framework’s resilience to quantum noise is also markedly enhanced because the trainable parameters reside entirely in the classical domain, insulating gradient updates from the stochasticity of quantum measurements and hardware imperfections. Moreover, the low-rank structure inherent in the TT representation further contributes to noise suppression by effectively averaging measurement noise across the TT-core gradients, a phenomenon theoretically linked to variance reduction. Our Neural Tangent Kernel (NTK) [3, 19, 25] analysis supports the improved trainability afforded by this structured parameterization.
@@
Our theoretical investigations confirm the enhanced representation capability, trainability, and generalization properties of TensorHyper-VQC. Experimentally, TensorHyper-VQC demonstrates superior performance and robustness across diverse applications, including quantum machine learning (e.g., quantum dot classification [20, 11]), combinatorial optimization (e.g., Max-Cut [49, 50]), and quantum simulation (e.g., LiH Hamiltonians [31, 13]). In all cases, TensorHyper-VQC outperforms conventional VQC and exhibits strong noise resilience across various noise models. These findings establish TensorHyper-VQC as a powerful and versatile computational framework that advances state-of-the-art variational quantum computing.
@@ 2.1 The TensorHyper-VQC Framework
Figure 2: The architecture of variational quantum computing. The classical input data are encoded as quantum states $|x_{1}\rangle$, $|x_{2}\rangle$, …, $|x_{U}\rangle$ via data encoding (e.g., angle or amplitude encoding). The shaded region denotes $L$ repeated parameterized quantum circuit block, consisting of trainable rotation gates $R_{X}(\alpha^{(l)}_{u})$, $R_{Y}(\beta^{(l)}_{u})$, $R_{Z}(\gamma^{(l)}_{u})$, and entangling layers across quantum channels. These blocks are stacked to increase circuit depth and expressive capacity. Finally, the circuit is measured in the Pauli-Z basis, yielding expectation values $\langle\sigma_{z}^{(1)}\rangle$, $\langle\sigma_{z}^{(2)}\rangle$, …, $\langle\sigma_{z}^{(U)}\rangle$ that serve as outputs for optimizaiton.
Throughout this work, we employ a single, consistent VQC architecture across all experiments and applications. The parameterized quantum circuit block, illustrated in Figure 2, is fixed in structure: it consists of alternating layers of single-qubit rotations $R_{X}(\alpha^{(l)}_{u})$, $R_{Y}(\beta^{(l)}_{u})$, $R_{Z}(\gamma^{(l)}_{u})$ and entangling gates across $U$ qubits, stacked to achieve the desired circuit depth $L$. By holding the VQC design constant, our results isolate the contribution of the TensorHyper-VQC parameterization strategy, rather than relying on task-specific ansatz engineering. This highlights that the observed gains in trainability, robustness, and scalability stem directly from the TT-based hypernetwork, demonstrating its versatility as a general-purpose parameter generator for VQC.
@@
Finally, a consistent TT hypernetwork architecture is adopted across all applications. The TT ranks and core dimensions are chosen to balance expressivity and generalization, while the same generator design is used across tasks. This uniformity underscores that the improvements in scalability, robustness, and trainability derive from the TT-guided parameterization principle itself rather than from task-specific quantum architectural search.
@@ 2.2 Theoretical Guarantees
$\mathcal{T}_{\rm tv}=\nabla_{\boldsymbol{\theta}}f_{\boldsymbol{\theta}}(\textbf{x})\nabla_{\boldsymbol{\theta}}f_{\boldsymbol{\theta}}(\textbf{x})^{\top}=\left(\frac{\partial f_{\boldsymbol{\theta}}(\textbf{x})}{\partial\hat{\textbf{w}}}\cdot\frac{\partial\hat{\textbf{w}}}{\partial\boldsymbol{\theta}}\right)\left(\frac{\partial f_{\boldsymbol{\theta}}(\textbf{x})}{\partial\hat{\textbf{w}}}\cdot\frac{\partial\hat{\textbf{w}}}{\partial\boldsymbol{\theta}}\right)^{\top}.$ (8)
@@
$\lambda_{\rm min}(\mathcal{T}_{\rm vqc})=\lambda_{\rm min}(\textbf{J}_{\hat{\textbf{w}}}\textbf{J}_{\hat{\textbf{w}}}^{\top}).$ (11)
@@
This shows that the effective NTK eigenvalue is not smaller and, under mild dispersion assumptions, can be strictly larger, thereby ensuring faster convergence and mitigating vanishing gradients. Accordingly, at epoch $t$, the optimization error satisfies:
@@
$\frac{\partial\hat{\mathcal{R}}}{\partial\mathcal{G}_{k}}=\left(\frac{\partial\hat{\mathcal{R}}}{\partial\hat{\textbf{w}}}+\boldsymbol{\tau}\right)\cdot\frac{\partial\hat{\textbf{w}}}{\partial\mathcal{G}_{k}}=\sum\limits_{u=1}^{3UL}\frac{\partial\mathcal{R}}{\partial\hat{w}_{u}}\cdot\frac{\partial\hat{w}_{u}}{\partial\mathcal{G}_{k}}+\sum\limits_{u=1}^{3UL}\tau_{u}\frac{\partial\hat{w}_{u}}{\partial\mathcal{G}_{k}}.$ (18)
@@ 2.4 Application to Quantum Machine Learning: Quantum Dot Classification
We first apply TensorHyper-VQC to a binary classification task on charge-stability diagrams from semiconductor quantum dots [44, 20, 14] to validate our theoretical framework. The objective is to classify diagrams into single-dot (Label 0) or double-dot (Label 1) systems, a crucial task for characterizing and controlling quantum devices. The experimental dataset comprises $50\times 50$-pixel images associated with quantum dot diagrams, including $2,000$ diagrams generated under realistic device-level noise conditions. We randomly partition the diagrams into $1,800$ training and $200$ test samples to evaluate the models’ generalization performance.
@@
The proposed TensorHyper-VQC architecture integrates a classical TT network as a parameter generator for VQC. The TT network is configured with input dimensions $[5,10,5,10]$, output dimensions $[4,2,3,9]$, and TT-ranks $[1,2,2,2,1]$, resulting in a compact representation with only $576$ trainable parameters (including bias). The TT network maps vectorized quantum-dot features to all rotation angles required by the VQC, thereby parameterizing $20$ qubits across $6$ variational layers (360 quantum-gate parameters). All models are trained using the cross-entropy loss [27] based on the Adam optimizer [22] with a fixed learning rate $3\times 10^{-2}$. We use a fixed random seed across all experiments to ensure reproducibility and fairness, and we initialize parameters from a normal distribution with mean 0 and variance 1.
To comprehensively benchmark performance, we compare TensorHyper-VQC against multiple baselines, including (i) a standard VQC with randomly initialized parameters, (ii) MLPHyper-VQC, where a multilayer perceptron replaces the TT network as the parameter generator, (iii) TTN+VQC, a hybrid model combining a TT encoder and VQC [35, 8], and (iv) ResNet18 [43] and ResNet50 [46] enhanced with low-rank adaptation (LoRA) [17, 26] for classical comparison. Furthermore, we conduct robustness analyses by injecting a variety of realistic noise models into the quantum circuits, including depolarizing, dephasing, two-Pauli, and readout noise, with error probabilities ranging from $0.1\%$ to $0.3\%$, simulating typical NISQ-era conditions.
To further assess resilience to noise, we incorporate standard barren-plateau mitigation and error-reduction techniques, including layer-wise training (LWT) [41], which progressively optimizes circuit depth to maintain gradient signal, and local cost function (LCF) design [7], which mitigates gradient vanishing by focusing on local observables. In addition, we employ standard quantum error mitigation approaches, including zero-noise extrapolation (ZNE) [15], readout error mitigation (REM) [42], and their combination (REM+ZNE). We evaluate the performance of all these strategies on both simulated and physical backends. Finally, we validate our approach on a 156-qubit IBM Quantum processor (Heron r2), demonstrating the deployability of TensorHyper-VQC on real hardware and its robustness to device-level imperfections, including decoherence, crosstalk, and measurement errors.
@@
Robustness Under Multi-Noise Conditions. We next evaluate the robustness under realistic quantum noise models [10, 38], including depolarizing, dephasing, two-Pauli, and readout noise channels, as illustrated in Fig. 3(c). Even as noise levels as high as $0.3\%$, TensorHyper-VQC maintains strong performance ($\geq 85\%$), significantly outperforming noise-agnostic VQC baselines. This robustness arises from the TT-based parameterization, which smooths the optimization landscape and enhances noise tolerance. Furthermore, as shown in Fig. 3(d), TenosrHyper-VQC exhibits graceful degradation when scaling from $20$ qubits down to $16$ or $12$ qubits, demonstrating robustness to hardware resource constraints while retaining competitive performance.
Comparison with Quantum Error Mitigation. We also compare TensorHyper-VQC with standard quantum error mitigation techniques, including ZNE, REM, and their combination as shown in Fig. 3(e). Remarkably, TensorHyper-VQC achieves a final accuracy of $94.0\%$, outperforming VQC enhanced with ZNE+REM ($55.7\%$) by nearly $40$ percentage points. This highlights that our tensor-structured parameterization intrinsically mitigates noise by reducing parameter variance and constraining quantum state evolution, often outperforming post-hoc mitigation methods that require additional circuit executions and calibration overhead.
@@
Table 1: Ablation study on $T$-rank selection under the TensorHyper-VQC’s configuration and noisy quantum setting described in Fig. (e).